Upon driving onto a freeway, a 1200kg car accelerates from a velocity of 15ms^(-1) to a velocity of 35ms^(-1). Calculate the net force that is acting on the car


1
Expert's answer
2021-08-12T06:22:12-0400

The net force that is acting on the car

"F=ma=m\frac{v_2-v_1}{t}"

For time "t=5\:\rm s" we get

"F=1200*\frac{35-15}{5}=4800\:\rm N"